# Improving thermal state preparation of Sachdev-Ye-Kitaev model with reinforcement learning on quantum hardware

### v1

@@ 4 Results
We employ a reinforcement learning (RL) agent-driven variational quantum thermal state algorithm introduced in Ref. [39]. For the dense SYK model, we consider the first circuit (PQC1, which calculates the entropy) to be constructed using strongly entangling layers [45]. The second circuit (PQC2 which calculates the expectation value) is built using the agent, as described also in the previous section. For the numerical simulations, we use Qulacs [46] implementation in a hybrid CPU and GPU interface [^4] CPU: AMD Rome and GPU: Nvidia Ampere A100.
Throughout the paper we use the gradient-free optimizer COBYLA [47] with $10^{3}$ iterations to optimize the parameters of the PQC1 and the PQC2. The RL framework is running for a maximum of $5\times 10^{3}$ episodes the RL environment interacts with the agent for a predefined number of steps. Furthermore, the agent is optimized by the ADAM optimzier [48] with $10^{-3}$ learning rate. Further details of the RL-agent can be found in Appendix C and the environment that contains the tensor-based state and the reward function is elaborated in Section 3.
@@
Overall, this post-processing step is crucial for refining our results and ensuring a more accurate representation of thermal states. In Appendix G we illustrate the impact of a wide range of the weights on the sampling of the best thermal state from all possible solutions proposed by the RL-agent.
@@ 4.1 Noiseless simulation
In Fig.3, we present the results for the dense SYK model simulations for $N=8,10$ Majorana fermions, the results are scaled up to $N=14$ in Fig. 4. Each simulation was carried out for five instances, and to benchmark the performance of the RL framework, the agent was retrained for each instance. In each panel of Fig.3 and Fig. 4, we report the free energy, the expectation value of the Hamiltonian, and the entropy. The solid line represents the mean of the exact values, while the green (yellow) region indicates the $1\sigma$ ($2\sigma$) deviation from the mean. Red dots correspond to the performance of the RL-agent parameterized circuit optimized using the free energy error as a reward signal (Eq. 4). Meanwhile, blue crosses depict the performance of the RL circuit using both the free energy error and the fidelity of the thermal state as reward signals (see Eq. 5). The error bars indicate one standard deviation for the given number of model instances. For $8\leq N\leq 12$, the fidelity achieved was in the range $0.80\leq F(\rho(\vec{\phi}))\leq 0.98$, whereas for $N=14$, the fidelity decreased below $0.80$ in the range $0.4<F(\rho(\vec{\phi}))\leq 0.80$.
@@ 4.2 Noise robustness on QPU
In this section, we study the robustness of the circuits selected by the RL-agent under realistic noisy conditions. The RL-agent is trained in a noiseless environment to optimize the parameterized circuits for minimizing the free energy. After training, the best circuit is selected according to how accurately it can find the free energy and then the circuit is executed on the IBM Eagle r3 processor to assess its practical performance. In the absence of any sophisticated error-mitigation techniques. See Appendix E for further information on the quantum hardware noise and limitations. The results were compared with exact classical calculations and experimental outcomes on the QPU for different temperatures.
@@ 5 Conclusion and future work
    A two order of magnitude reduction in the number of CNOT gates for systems $N\geq 12$ compared to traditional methods like first-order Trotterization.
@@
This approach represents a significant advancement in RL-based quantum circuit optimization, with potential applications extending beyond the SYK model to broader quantum simulations and quantum gravity studies on near-term quantum hardware.
@@
    Scaling to larger systems: Investigate the scalability of this approach for SYK models with $N>14$ Majorana fermions, advancing the limits of current quantum simulations. This can be achieved by leveraging the recently proposed gadget reinforcement learning [54], which enables the learning of composite gate sets from smaller Majorana fermionic SYK models and subsequently applying them to solve models with a larger number of Majorana fermions efficiently.
@@
    Out-of-time-order correlations: Apply this method to accurately compute out-of-time-order correlations (OTOCs) for the SYK model, which are crucial for studying quantum chaos and holographic properties.
@@
    Exploring reward engineering: Explore other forms of reward function to benchmark the performance of the CNN agent. Such as we can encode the information corresponding to the properties of the thermal state (e.g., the entropy, purity, etc.) to improve the thermal state preparation further.
@@
    Quantum information-theoretic analysis Recent quantum informatic research suggests analyzing RL-agent-produced circuits can significantly enhance agent performance [55]. By applying quantum information-theoretic principles to circuit analysis, we can optimize circuit architectures, identify performance bottlenecks, and refine the RL framework’s efficiency.
@@
    Advanced error mitigation: Implement sophisticated error mitigation techniques such as [56, 50, 53, 57, 51] to improve the accuracy of results in noisy quantum environments, particularly for free energy calculations.
@@ 6 Acknowledgment
The author expresses my gratitude to Sara Doelman for her insightful discussions on convolutional neural networks and her encouragement to publish these results. The author also thanks Raghav Govind Jha for the conversations that inspired this paper. Additionally, the author is grateful to Guilherme Ilário Correr, Aritra Sarkar, Abhishek Sadhu, Jarosław Adam Miszczak, and Stefano Mangini for their valuable suggestions and comments that improved the quality of this article. Finally, The author wish to acknowledge CSC–IT Center for Science, Finland, for computational resources.
Appendix A Reinforcement learning framework
@@
Appendix B Tensor-based encoding of quantum circuits
We adopt a binary encoding scheme, as introduced in [28, 29], to capture the structure of PQC2, focusing on the order and arrangement of gates to provide the agent with a complete circuit description. Unlike previous approaches, which flattened the 3D encoding into one dimension for use in feedforward neural networks, our method leverages the full spatial structure of the encoding by employing a 3D convolutional neural network (CNN). This allows us to extract richer spatial features from the encoded representation as shown in the Appendix C. .
@@ Appendix C CNN outperforms FNN
Appendix D Resource quantization
@@
Appendix E The IBM Eagle r3 processor
@@
Appendix F The agent and environment hyperparameters
@@
    Final $\gamma$: The discount factor determining the importance of future rewards compared to immediate rewards.
@@
Appendix G The art of sampling best circuits
@@
where $\Delta F$, $\Delta\langle H\rangle_{\beta}$, and $\Delta S$ represent errors in free energy, Hamiltonian expectation, and entropy, respectively. Weights $w_{a}$ and $w_{b}$ adjusts the relative importance of these errors. The best-performing circuit is selected by:
  1. 1.
    Evaluating each proposed circuit using $\varepsilon$.
  2. 2.
    Selecting the circuit that minimizes $\varepsilon$.
Figures 10 and 11 illustrate the significant impact of weight selection on circuit sampling. These results emphasize the critical nature of careful postprocessing to achieve an optimal balance between free energy, Hamiltonian expectation value, and entropy of the thermal state. The findings underscore the importance of thoughtful parameter selection in the filtering process, as it directly influences the quality of the sampled circuits. This approach allows for fine-tuning the selection criteria to meet specific performance requirements in quantum circuit design and optimization.

### v2

@@ 4 Results
We implement a reinforcement learning (RL)-driven hybrid quantum-classical thermal state preparation algorithm (see Ref. [42] for the details on the variational algorithm). The entropy-estimation circuit $\mathrm{PQC}_{1}$ employs entangling layers with cyclic connectivity, while $\mathrm{PQC}_{2}$, which calculates the expectation value, is built using the RL-agent. Numerical simulations leverage the Qulacs quantum simulator [48] in a heterogeneous architecture [^5] Hybrid computation: 64-core AMD EPYC CPUs and NVIDIA A100 GPUs via CUDA-quantum integration [amd2025heterogeneous].
Throughout the paper we use the gradient-free optimizer COBYLA [49] with $10^{3}$ iterations to optimize the parameters of the PQC1 and the PQC2. The RL framework is running for a maximum of $5\times 10^{3}$ episodes the RL environment interacts with the agent for a predefined number of steps. Furthermore, the agent is optimized by the ADAM optimzier [50] with $10^{-3}$ learning rate. Further details of the RL-agent can be found in Appendix F and the environment that contains the tensor-based state and the reward function is elaborated in Section 3.
@@
Overall, this post-processing step is crucial for refining our results and ensuring a more accurate representation of thermal states. In Appendix K, we illustrate the impact of a wide range of weights on the sampling of the best thermal state from all possible solutions proposed by the RL-agent.
@@ 4.1 Noiseless simulation
Figures 3 and 4 demonstrate our framework’s performance for the dense SYK model across system sizes $N=8,10$ (Fig. 3) and $N=14$ (Fig. 4). Five disorder realizations were simulated per system size, with agent retraining per instance to ensure robust benchmarking. Each panel quantifies three thermodynamic quantities: the free energy $F=-\beta^{-1}\ln Z$, energy expectation $\langle H\rangle$, and entropy $S=\beta(\langle H\rangle-F)$. Shaded regions (green: $1\sigma$, yellow: $2\sigma$) represent statistical deviations from exact diagonalization results (solid lines). To this end, two reward strategies are compared:
  • Red markers: Free energy minimization via reward. 4.
  • Blue markers: Combined free energy and state fidelity rewards in Eq. 18.
Error bars reflect standard deviations across disorder realizations. The fidelity $F(\rho(\boldsymbol{\phi}))$ remains high (0.800.800.800.800.980.980.980.98) for $8\leq N\leq 12$, but degrades substantially (0.40.40.40.40.800.800.800.80) at $N=14$, highlighting the challenge of maintaining state quality in larger systems.
@@ 4.2 Noise robustness on QPU
In this section, we study the robustness of the circuits selected by the RL-agent under realistic noisy conditions. The RL-agent is trained in a noiseless environment to optimize the parameterized circuits for minimizing the free energy. After training, the best circuit is selected according to how accurately it can find the free energy and then the circuit is executed on the IBM Eagle r3 processor to assess its practical performance. In the absence of any sophisticated error-mitigation techniques. See Appendix H for further information on the quantum hardware noise and limitations. The results were compared with exact classical calculations and experimental outcomes on the QPU for different temperatures.
@@ 5 Conclusion and future work
    A two-order-of-magnitude reduction in the number of CNOT gates for systems $N\geq 12$ compared to traditional methods like first-order Trotterization.
@@
While our current implementation does not match the system sizes achievable in classical simulations of the SYK model (up to $N\sim 60$ [9]), this work demonstrates important progress in RL-optimized quantum circuits. The methodology developed here has broader implications for quantum simulation, particularly for studying quantum gravity phenomena on near-term hardware.
@@
    Scaling to larger systems by gadgets: We propose to investigate the scalability of our approach for SYK models with $N>14$ Majorana fermions, pushing the boundaries of current quantum simulations. This can be achieved by leveraging the recently proposed gadget reinforcement learning [56], which enables the learning of composite gate sets from smaller Majorana fermionic SYK models. These learned gate sets can then be efficiently applied to solve models with a larger number of Majorana fermions, enhancing scalability and providing insights into the structure of the quantum circuits, including identifying optimal gate sequences, circuit compression, and potential generalization to other quantum systems or problems.
@@
    Out-of-time-order correlations: Apply this method to accurately compute out-of-time-order correlations (OTOCs) for the SYK model [28], which are crucial for studying quantum chaos and holographic properties.
@@
    Exploring reward engineering: Explore other forms of reward function to benchmark the performance of the CNN agent. Such as, we can encode the information corresponding to the properties of the thermal state (e.g., the entropy, purity, etc.) to improve the thermal state preparation further. For instance, extending Eq. 18 to include weighted entropy optimization:
    $r=\begin{cases}5,&\text{if }F_{t}\leq\zeta_{F}\land\textrm{Fid}(\rho({\vec{\phi}}))\geq\zeta_{\textrm{Fid}}\land\textrm{Ent.}_{\text{term}}\leq\zeta_{\text{Ent.}},\\ -5,&\text{if }\textrm{Fid}(\rho({\vec{\phi}}))<\zeta_{\textrm{Fid}}\land\text{step no.}=D_{\text{max}},\\ aE_{\text{term}}+b\textrm{Fid}_{\text{term}}+c\textrm{Ent.}_{\text{term}},&\text{otherwise},\end{cases}$ (8)
    where $c$ controls entropy weighting, could better align state preparation with specific observables.
@@
    Transfer learning and policy generalization: In this study, the RL-agent is trained from scratch for each inverse temperature $\beta$ and for different values of $N$. However, one could implement transfer learning strategies [57] to adapt the policy learned for a specific $N$ and reuse it across different inverse temperatures.
@@
    Quantum information-theoretic analysis: Recent quantum informatic research suggests analyzing RL-agent-produced circuits can significantly enhance agent performance [58]. By applying quantum information-theoretic principles to circuit analysis, we can optimize circuit architectures, identify performance bottlenecks, and refine the RL framework’s efficiency.
@@
    Advanced error mitigation: Implement sophisticated error mitigation techniques such as [59, 52, 55, 60, 53] to improve the accuracy of results in noisy quantum environments, particularly for free energy calculations for larger $N$.
  7. 7.
@@ 6 Acknowledgment
The author expresses my gratitude to Sara Doelman for her insightful discussions on convolutional neural networks and her encouragement to publish these results. The author also thanks Raghav Govind Jha for the conversations that inspired this paper. Additionally, the author is grateful to Guilherme Ilário Correr, Aritra Sarkar, Abhishek Sadhu, Jarosław Adam Miszczak, and Stefano Mangini for their valuable suggestions and comments that improved the quality of this article. Finally, The author wishes to acknowledge CSC–IT Center for Science, Finland, for computational resources. The author acknowledges funding from the Research Council of Finland through the Finnish Quantum Flagship project 358878 (UH).
Appendix A In depth analysis of CNOT scaling
In Fig. 7, we demonstrate that the number of CNOT gates scales polynomially with qubit count, approximately following the cubic trend
$y=2.06N_{q}^{3}-27.73N_{q}^{2}+123.54N_{q}-168.71N_{q},$ (9)
where $N_{q}$ is the number of qubits. This represents a substantial advantage over the exponential scaling of real-time dynamics
$y=14.839e^{0.890N_{q}}-379.902.$ (10)
Figure 7: Scaling of CNOT gate counts with qubit number for our method (polynomial fit) versus real-time dynamics (exponential fit). The cubic scaling of our approach ($y=2.06N_{q}^{3}-27.73N_{q}^{2}+123.54N_{q}-168.71$) offers a substantial advantage over the exponential growth observed in real-time dynamics ($y=14.839e^{0.890N_{q}}-379.902$), particularly for larger systems. Fitted curves and data points are shown for 3–10 qubits, highlighting the divergence in gate requirements as system size increases. Analytical Trotterization results for 10 qubits are also indicated for comparison.
This scaling difference becomes crucial for larger systems: our method maintains $\mathcal{O}(N_{q}^{3})$ CNOT growth with qubit number, while real-time dynamics exhibits $\mathcal{O}(e^{0.890N_{q}})$ scaling. The fitted curves reveal comparable CNOT counts for small systems: $8$ (ours) vs. $30$ (real-time) for 3-qubit systems and $13$ vs. $110$ for 4-qubit systems. However, the counts diverge dramatically as $N$ increases. This polynomial growth pattern suggests feasibility for larger SYK simulations on near-term hardware. Notably, the exponential scaling equation overestimates gate counts at higher qubit numbers: for 10-qubit systems, it predicts $\sim\!230000$ gates, while analytical Trotterization yields $19984$ CNOTs per layer [28]. With two layers ($39968$ total), this remains significantly below the exponential prediction. This discrepancy arises from limited data points during curve fitting.
The confidence intervals (CI) visualized in the green shaded regions were computed using the delta method [61] for uncertainty propagation in nonlinear models. For the exponential model $f(x)=a\cdot e^{bx}+c$, the Jacobian matrix $J$ was constructed containing partial derivatives to each parameter ($\partial f/\partial a=e^{bx}$, $\partial f/\partial b=ax\cdot e^{bx}$, and $\partial f/\partial c=1$). The variance of the prediction at each point $x_{i}$ was then calculated as $\sigma_{y_{i}}^{2}=J_{i}^{T}\Sigma J_{i}$, where $\Sigma$ is the parameter covariance matrix obtained from the curve-fitting procedure. For a confidence level $\alpha$ (0.05 for 95% confidence), the intervals were computed as $\hat{y}i\pm t{n-p,1-\alpha/2}\cdot\sqrt{\sigma_{y_{i}}^{2}}$, where $t_{n-p,1-\alpha/2}$ is the critical value from Student’s t-distribution [62] with $n-p$ degrees of freedom ($n$ being the number of data points and $p$ the number of parameters). The widening confidence bands at higher qubit counts reflect the increasing uncertainty in predictions as we extrapolate beyond the measured data range, with the exponential model showing particularly large uncertainties due to the compounding effect of parameter uncertainty in the exponent.
Appendix B Resource requirements and scalability of entropy measurement from PQC1
Accurate entropy measurement in quantum systems necessitates a careful balance between precision and resource efficiency. Traditional approaches such as quantum state tomography reconstruct the full density matrix through repeated state preparation and measurement but suffer from exponential scaling, requiring $O(3^{n})$ measurements for $n$ qubits. This limits practical applications to systems with $n\gtrsim 10$ qubits. Modern methodologies achieve polynomial efficiency under specific constraints:
Variational quantum diagonalization employs parameterized quantum circuits optimized to estimate von Neumann entropy with $O(\mathrm{poly}(n))$ measurements [63, 31]. While measurement-efficient, these methods incur classical optimization overhead that scales with circuit depth. Classical shadows leverage randomized single-qubit measurements to estimate the second Rényi entropy with $O(\log n)$ measurements per state copy [64], though they provide limited insight into mixed-state von Neumann entropy. Hybrid approaches combining machine learning methods with local measurements enable entanglement entropy prediction for systems up to $\sim 100$ qubits [64], albeit with assumptions about spatial entanglement structure.
Scalability depends critically on state purity. For low-rank states (e.g., pure or low-temperature thermal states), variational and shadow-based methods exhibit polynomial scaling, demonstrated experimentally for $n\leq 50$ qubits [65]. In contrast, high-entropy full-rank systems face fundamental $O(2^{n})$ measurement bottlenecks due to Hilbert space growth. All techniques require $O(1/\epsilon^{2})$ measurements for precision $\epsilon$, with error mitigation strategies [53, 66] becoming essential on noisy hardware. Recent advances include reinforcement learning-optimized protocols that adaptively minimize measurement rounds [31] and symmetry-aware shadow estimation reducing sampling costs by 60% for states with conserved quantities [65]. While polynomial-scaling methods enable practical entropy estimation for intermediate-scale quantum devices ($n\sim 100$), challenges persist for highly mixed states and systems with non-Markovian noise. Hybrid quantum-classical algorithms leveraging problem-specific structure currently offer the most viable path forward, though fundamental limitations remain for generic high-entropy systems.
Appendix C Elaboration on variational thermal state preparation
The task of variational thermalization is: given a Hamiltonian $H$ and a target temperature $T=1/\beta$, we need to prepare $\rho_{\beta}$ (the thermal state), which is expressed by
$\rho_{\beta}=\frac{e^{-\beta H}}{\operatorname{tr}(e^{-\beta H})},$ (11)
where the partition function $Z_{\beta}=\operatorname{tr}(e^{-\beta H})$. To implement this in a variational quantum algorithm, we consider a quantum ansatz state $\rho_{\theta\phi}$ and define the relative entropy between the ansatz state and the thermal state as
$D(\rho_{\theta\phi}||\rho_{\beta})=-S(\rho_{\theta\phi})-\operatorname{tr}(\rho_{\theta\phi}\log\rho_{\beta}),$ (12)
which can be rewritten using Eq. 11 as
$D(\rho_{\theta\phi}||\rho_{\beta})=-S(\rho_{\theta\phi})+\beta\operatorname{tr}(\rho_{\theta\phi}H)+\log Z_{\beta}.$ (13)
This is known as the relative quantum entropy. Using the free energy definition $F(\cdot)\equiv\operatorname{tr}((\cdot)H)-\frac{S(\cdot)}{\beta}$, we can rewrite this as
$D(\rho_{\theta\phi}||\rho_{\beta})=\beta F(\rho_{\theta\phi})-\beta F(\rho_{\beta}).$ (14)
Thus, $\rho_{\theta\phi}$ represents the thermal state $\rho_{\beta}$ iff $F(\rho_{\theta\phi})=F(\rho_{\beta})$ and $\rho_{\theta\phi}=\rho_{\beta}$, i.e.,
$D(\rho_{\theta\phi}||\rho_{\beta})=0.$ (15)
Therefore, for a variational thermal state ansatz, minimizing the free energy
$C(\rho_{\theta\phi})=\beta F(\rho_{\theta\phi})=\beta\operatorname{tr}(\rho_{\theta\phi}H)-S(\rho_{\theta\phi})$ (16)
yields optimal parameters where $\rho_{\theta\phi}=\rho_{\beta}$.
In a state-of-the-art [67] implementation, entropy calculation occurs classically while energy expectation is on a quantum computer. To compute both on a quantum computer [42], as we do in our case:
  • $\rho_{\theta}$ is generated via a parameterized quantum circuit $\mathrm{PQC}_{1}(\theta)$. We use $\rho_{\theta}$ to calculate the entropy $S(\rho_{\theta\phi})$.
  • $\rho_{\theta\phi}$ is obtained by initialized another parameterized quantum circuit $\mathrm{PQC}_{2}(\phi)$ using $\rho_{\theta}$. Then $\rho_{\theta\phi}$ is used to calculate the Hamiltonian expectation value $E(\rho_{\theta\phi})$.
  • $E(\rho_{\theta\phi})$ and $S(\rho_{\theta\phi})$ compute the free energy $\beta F(\rho_{\theta\phi})$ for specific $\beta$.
Equations 15 and 16 show that $D(\rho_{\theta\phi}||\rho_{\beta})=0$ implies $\rho_{\theta\phi}=\rho_{\theta}=\rho_{\beta}$. Hence, during optimization $\rho_{\theta}\neq\rho_{\theta\phi}$ but when the loss $C(\rho_{\theta\phi})$ is minimized we get $\rho_{\theta}=\rho_{\theta\phi}=\rho_{\beta}$.
Appendix D Reinforcement learning framework
@@
Appendix E Tensor-based encoding of quantum circuits
We adopt a binary encoding scheme, as introduced in [30, 31], to capture the structure of PQC2, focusing on the order and arrangement of gates to provide the agent with a complete circuit description. Unlike previous approaches, which flattened the 3D encoding into one dimension for use in feedforward neural networks, our method leverages the full spatial structure of the encoding by employing a 3D convolutional neural network (CNN). This allows us to extract richer spatial features from the encoded representation as shown in the Appendix F. .
@@ Appendix F CNN outperforms FNN
Method Error Parameters Depth CNOT
CRLQAS [30] $2.6\times 10^{-6}$ 29 22 11
KANQAS [44] $1.1\times 10^{-4}$ $\mathbf{15}$ 18 11
Ours $\mathbf{4.42\times 10^{-7}}$ 19 $\mathbf{16}$ $\mathbf{7}$
Table 3: Comparison of quantum circuit metrics.
Fig. 9 not only demonstrates the clear advantage of using CNNs over FNNs within the RL framework but also highlights their advantage over state-of-the-art quantum architecture search algorithms, such as KANQAS [44] and CRLQAS [30]. Specifically, the proposed approach achieves higher accuracy in finding the ground state of chemical Hamiltonians while utilizing smaller quantum circuits. A detailed comparison of these results is provided in Tab. 3.
Appendix G Training time
In Tab. 4 we discuss the noiseless/noisy simulation time by various SYK Hamiltonians. All the simulation is done under a computational budget of 48 hours within which the RL-framework completed approximately $5\times 10^{3}$ episodes.
@@ Appendix G Training time
Appendix H The IBM Eagle r3 processor
@@
Appendix I The agent and environment hyperparameters
@@
    Final $\gamma$: The discount factor determines the importance of future rewards compared to immediate rewards.
@@
Appendix J Reward function analysis
In this section, we discuss the effect of weight corresponding to free energy ($a$ and the weight corresponding to Hamiltonian expectation ($b$) on the reward function
$r=\begin{cases}5,&\text{if }F_{t}(\beta=\text{const.},\vec{\theta},\vec{\phi}){\leq}\zeta_{F}\;\text{and,}\;\textrm{Fid}(\rho({\vec{\phi}}))\geq\zeta_{\textrm{Fid.}}\\ -5,&\text{if }\textrm{Fid}(\rho({\vec{\phi}}))<\zeta_{\textrm{Fid}}\;\text{and,}\;\text{step no.}=D_{\text{max}},\\ a\times E_{\text{term}}+b\times\textrm{Fid}_{\text{term}},&\text{otherwise},\end{cases}$ (18)
The analysis in Tab. 6 reveals a nuanced relationship between the precision of free energy estimation and the focus on expectation values. As shown in Table 6, increasing the weight $a$ generally leads to improved precision in free energy estimation, with the lowest error of $2.15\times 10^{-7}$ observed at $a=0.8$ and $b=0.2$.
a b Free energy error Hamiltonian expectation error
0.0 1.0 $5.14\times 10^{-6}$ $7.84\times 10^{-4}$
0.2 0.8 $1.64\times 10^{-6}$ $2.67\times 10^{-3}$
0.4 0.6 $2.17\times 10^{-6}$ $8.35\times 10^{-4}$
0.6 0.4 $1.81\times 10^{-6}$ $2.73\times 10^{-3}$
0.8 0.2 $2.15\times 10^{-7}$ $2.53\times 10^{-3}$
1.0 0.0 $5.14\times 10^{-6}$ $7.84\times 10^{-4}$
Table 6: Free energy and expectation value errors for different reward function weight configurations. Where $a$ = weight corresponding to free energy, $b$ = weight corresponding to Hamiltonian expectation value.
However, this trend does not continue indefinitely. When $a$ is set to 1.0, effectively eliminating the focus on the Hamiltonian expectation value ($b=0.0$), we observe a decrease in precision, with the free energy error increasing to $5.14\times 10^{-6}$. This suggests that maintaining some weight on the expectation value is crucial for optimal performance.
These findings highlight the importance of balancing the emphasis between free energy and expectation value in the reward function. While prioritizing free energy estimation (higher $a$ values) generally improves precision, completely neglecting the expectation value component can lead to suboptimal results. Future work could explore finer-grained weight configurations to determine the optimal balance for specific applications.
Appendix K The art of sampling best circuits
@@
where $\Delta F$, $\Delta\langle H\rangle_{\beta}$, and $\Delta S$ represent errors in free energy, Hamiltonian expectation, and entropy, respectively. Weights $w_{a}$ and $w_{b}$ adjust the relative importance of these errors. The best-performing circuit is selected by: (1) Evaluating each proposed circuit using $\varepsilon$. (2) Selecting the circuit that minimizes $\varepsilon$.
Figures 11 and 12 illustrate the significant impact of weight selection on circuit sampling. These results emphasize the critical nature of careful postprocessing to achieve an optimal balance between free energy, Hamiltonian expectation value, and entropy of the thermal state. The findings underscore the importance of thoughtful parameter selection in the filtering process, as it directly influences the quality of the sampled circuits. This approach allows for fine-tuning the selection criteria to meet specific performance requirements in quantum circuit design and optimization.